X Industries has current assets equal to $9 million.

The company's current ratio is 2.0, and its quick ratio is 1.3.

What is the firm's level of current liabilities? What is the firm's level of inventories? Do not round intermediate calculations. Round your answers to the nearest dollar.
Answer #1

Current ratio=Current assets/current liabilities

current liabilities=(9/2)

=$4.5 million(or $4,500,000)

Quick ratio=(Current assets-inventory)/current liabilities

1.3=(9-inventory)/4.5

(1.3*4.5)=9-inventory

inventory=9-5.85

=$3.15 million=$3,150,000